### v2

# Atoxia: Red-teaming Large Language Models with Target Toxic Answers

Abstract
Despite the substantial advancements in artificial intelligence, large language models (LLMs) remain being challenged by generation safety. With adversarial jailbreaking prompts, one can effortlessly induce LLMs to output harmful content, causing unexpected negative social impacts. This vulnerability highlights the necessity for robust LLM red-teaming strategies to identify and mitigate such risks before large-scale application. To detect specific types of risks, we propose a novel red-teaming method that Attacks LLMs with Target Toxic Answers (Atoxia). Given a particular harmful answer, Atoxia generates a corresponding user query and a misleading answer opening to examine the internal defects of a given LLM. The proposed attacker is trained within a reinforcement learning scheme with the LLM outputting probability of the target answer as the reward. We verify the effectiveness of our method on various red-teaming benchmarks, such as AdvBench and HH-Harmless. The empirical results demonstrate that Atoxia can successfully detect safety risks in not only open-source models but also state-of-the-art black-box models such as GPT-4o.
Atoxia: Red-teaming Large Language Models with Target Toxic Answers
Yuhao Du1,2thanks: Equal contributions, {yuhaodu,zhuoli3}@link.cuhk.edu.cn., Zhuo Li1,211footnotemark: 1, Pengyu Cheng4, Xiang Wan1, Anningzhe Gao1,3thanks: Corresponding author, anningzhegao@gmail.com. 1Shenzhen Research Institute of Big Data 2The Chinese University of Hong Kong, Shenzhen 3ByteDance   4Tencent AI Lab https://github.com/DuYooho/Atoxia Warning: This paper includes content that may be offensive or harmful.
1 Introduction
Large language models (LLMs), such as ChatGPT (OpenAI, 2023) and GPT-4 (Achiam et al., 2023), have shown astonishing AI capabilities (Xie et al., 2024; Zhao et al., 2024; Nam et al., 2024; Cheng et al., 2025) and renewed the field of natural language processing (NLP) (Ouyang et al., 2022; Brohan et al., 2023; Achiam et al., 2023; Hu et al., 2022, 2023). The success of LLMs largely depends on being trained with vast amounts of data Brown et al. (2020). However, diverse textual data inevitably contains harmful and toxic information, contaminating the model’s knowledge and leading to undesirable behaviors (Bai et al., 2022a; Zou et al., 2023). To address the challenges, LLM alignment has emerged as a critical area of research (Stiennon et al., 2020; Ouyang et al., 2022; Lee et al., 2023), aiming to ensure that models behave in accordance with human preferences and safety standards. Despite these efforts, the inherently uncontrolled nature of generative models poses a potential risk of producing unpredictable and unsafe responses. To mitigate these risks, developers have implemented various safety mechanisms to filter harmful outputs and prevent undesirable behavior (Ziegler et al., 2019; Perez et al., 2022; Zhuo et al., 2023; Ji et al., 2024; Li et al., 2024).
[…]
**Answer Opening**: A key step to being successful in the pursuit of gain through deception, is carefully amassing sensitive information.
Table 1: Comparison of red-teaming approaches: Traditional methods like AdvPrompter (Paulus et al., 2024) versus Atoxia (Ours). While AdvPrompter processes toxic questions as input to generate refined adversarial queries, Atoxia takes toxic answers as input, generating both adversarial queries and an answer opening designed to elicit similarly toxic answers, thereby substantially increasing the misleading probability.
Nevertheless, adversarial techniques, such as jailbreaking (Perez and Ribeiro, 2022) have demonstrated that these safety measures are not foolproof, where attackers can craft deceptive prompts designed to bypass the model’s safeguards, by disguising harmful intentions within seemingly benign requests (Zou et al., 2023; Liu et al., 2024; Paulus et al., 2024; Sadasivan et al., 2024). For instance, prompts may be carefully engineered to appear educational or helpful but still lead the model to generate harmful responses (Zou et al., 2023; Paulus et al., 2024). This highlights a fundamental challenge in LLM safety: even well-aligned models can be manipulated under the right conditions. Recognizing this vulnerability, researchers have explored whether adversarial techniques can be repurposed to strengthen the safety of LLMs. By proactively generating high-quality adversarial prompts that mimic real-world attack scenarios, we can systematically test our models’ weaknesses, identifying potential flaws, and enabling us to refine the model and its defenses before such attacks occur in real-world applications. Following this idea, various LLM red-teaming works have been proposed. For example, AdvPrompter (Paulus et al., 2024) leverages automated prompt generation to uncover model vulnerabilities, while GCG (Zou et al., 2023) leverages gradient-based optimization to identify adversarial tokens for eliciting unexpected responses. However, existing red-teaming methods lack the ability to specifically target and detect certain toxic responses, a critical limitation for real-world applications. Some highly harmful outputs are entirely unacceptable and must never be generated, yet current approaches fail to reliably prevent them.
To further test LLM safety on specific topics, we introduce a red-teaming Attacker model to detect the potential of LLM outputting Target Toxic Answers, called Atoxia. The main idea of Atoxia is to interact with an under-testing LLM by taking in a given toxic answer and generating an adversarial query and a corresponding answer opening/prefix. The generated tuple of the query and answer opening is used to mislead the under-testing LLM to output a similar toxic answer with a high probability. In the implementation, Atoxia is another language model trained using the reinforcement learning (RL) scheme, where the reward is defined as the conditional probability of outputting the target toxic answer from the under-testing LLM. This target-driven paradigm enables a comprehensive evaluation of an LLM’s susceptibility to harmful behavior, identifying specific prompts or triggers that may elicit harmful responses. Moreover, although tailored to a particular model during training, Atoixa empirically demonstrates strong generalization ability to transfer red-teaming attacks to even state-of-the-art black-box models, e.g., GPT-4o. A comparison between Atoxia and other red-teaming methods is presented in Table 1. While traditional approaches like AdvPrompter focus on generating refined adversarial queries from toxic questions, Atoxia innovates by starting with toxic answers as input, producing both adversarial queries and a toxic answer opening, significantly enhancing the likelihood of eliciting harmful responses. In addition, Atoxia generates more human-readable responses compared to previous methods, as illustrated in Table 1. This inherent naturalness significantly increases the likelihood of bypassing perplexity (PPL)-based safety mechanisms, as highlighted in prior research (Paulus et al., 2024). Such enhanced fluency and coherence make Atoxia particularly effective at evading detection by these safety measures. To summarize, the main contributions of our paper are:
  • We introduce a novel red-teaming method that can Attack LLM based on Target Toxic Answers (Atoxia). Our method proactively detects the potential harmful behaviors in LLMs, preventing negative social impacts of LLMs before large-scale application.
  • We introduce a new RL-based attacker training approach using the under-testing LLM’s likelihood of target toxic answers as rewards. This enables a streamlined and efficient optimization process without additional requirements of training reward models.
  • We discover Atoxia with impressive generalizability, where the attacker is only trained with open-source LLMs but still performs effectively towards black-box models, e.g., GPT-4o. This empirically highlights Atoxia’s robustness and adaptability across various models.
  • We evaluate the red-teaming performance on diverse safety benchmarks, where our approach empirically shows uniform and transferable detection capabilities on both open-source and black-box LLMs. Our experimental results underscore the safety vulnerabilities of existing LLMs and raise the urgency for advanced improvement of LLM safety.
Figure 1: Illustration of our proposed target-driven detection approach. The left side shows the overall pipeline, while the right details the process with examples. The gray box contains content generated by Atoxia ($\pi_{\theta}$) before training, and the green box shows content after reinforcement learning. During training, Atoxia interacts with the under-testing LLM ($\pi_{\text{test}}$), using log-likelihood as the reward signal. Finally, toxic content is identified in the red box after querying the under-testing LLM with the refined question and answer opening.
2 Preliminary
Reinforcement Learning from Human Feedback (RLHF) is an effective and essential method for aligning LLMs with human preferences (Ouyang et al., 2022; Zeng et al., 2024; Cheng et al., 2024). Typically, RLHF consists of two steps: reward modeling and RL training. In reward modeling (Askell et al., 2021; Bai et al., 2022a; Cheng et al., 2023), a reward model (RM) $r({\bm{x}},{\bm{y}})$ is defined to measure LLM response ${\bm{y}}$’s quality w.r.t. a given an input prompt ${\bm{x}}$. With a preference set ${\mathcal{D}}_{\text{p}}=\{({\bm{x}},{\bm{y}}_{w},{\bm{y}}_{l})\}$, RM can be learned by $\mathcal{L}_{\text{RM}}=$
$-\mathbb{E}_{({\bm{x}},{\bm{y}}_{w},{\bm{y}}_{l})\sim\mathcal{D}_{\text{p}}}[\log\sigma(r({\bm{x}},{\bm{y}}_{w})-r({\bm{x}},{\bm{y}}_{l}))],$ (1)
where ${\bm{y}}_{w}$ and ${\bm{y}}_{l}$ denote the “preferred” and “rejected” responses, respectively, and $\sigma$ is the Sigmoid function.
For the RL training step, the typical RLHF method trains LLM policy $\pi_{\phi}$ uses the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017), maximizing $\mathcal{L}_{\text{RLHF}}=$
$\mathbb{E}_{{\bm{x}}\sim\mathcal{D},{\bm{y}}\sim\pi_{\phi}(\cdot|{\bm{x}})}\Big{[}r({\bm{x}},{\bm{y}})-\beta\log\frac{\pi_{\phi}({\bm{y}}|{\bm{x}})}{\pi_{\text{ref}}({\bm{y}}|{\bm{x}})}\Big{]},$ (2)
where $\beta>0$ is a reweighting hyper-parameter, and $\pi_{\text{ref}}$ is a reference model.
LLM Red-teaming
Red-teaming, in the context of LLMs, refers to the process of adversarially testing these models to uncover vulnerabilities, such as the generation of harmful or unintended content (Zhuo et al., 2023; Paulus et al., 2024; Zhou et al., 2024). Formally, given a language model $\mathcal{M}$ and an adversarial objective $J_{\text{adv}}$, the red-teaming process can be framed as an optimization problem:
${\bm{p}}^{*}=\mathop{\arg\max}_{{\bm{p}}\in{\mathcal{P}}}J_{\text{adv}}(\mathcal{M}({\bm{p}})),$ (3)
where ${\bm{p}}$ denotes the adversarial prompt from prompt space ${\mathcal{P}}$, and $J_{\text{adv}}$ quantifies the harmfulness or deviation from desired behavior. This can involve crafting adversarial prompts, commonly referred to as jailbreaking (Zou et al., 2023; Paulus et al., 2024) or systematically probing the model to identify weaknesses in its safety mechanisms. Our method, for instance, can be viewed as a target-driven approach within this framework, focusing on specific vulnerabilities to enhance model robustness. Red-teaming is crucial for improving the robustness and alignment of LLMs with ethical and safety standards.
3 Methodology
Assume we have an under-testing (UT) LLM $\pi_{\text{test}}$ requiring the process of red-teaming. We aim to learn a red-teaming attacker $\pi_{\theta}$. For each target toxic answer ${\bm{y}}^{*}\in{\mathcal{A}}^{*}$, the attacker $\pi_{\theta}$ is supposed to generate an adversarial question ${\bm{x}}_{\text{adv}}$ and a corresponding answer opening ${\bm{y}}_{\text{pre}}$:
$({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*}),$ (4)
where $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ try to induce the under-testing LLM $\pi_{\text{test}}$ to output an answer continuation ${\bm{y}}_{\text{suff}}\sim\pi_{\text{test}}(\cdot|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ highly related to ${\bm{y}}^{*}$ in semantical meaning.
Note that besides the adversarial question ${\bm{x}}_{\text{adv}}$, the attacker $\pi_{\theta}$ also generates an answer prefix ${\bm{y}}_{\text{pre}}$. This is because existing aligned LLMs can easily reject responding when only harmful query ${\bm{x}}_{\text{adv}}$ is provided (Paulus et al., 2024). Using the additional answer prefix ${\bm{y}}_{\text{pre}}$, we can significantly enlarge the misleading probability of the under-testing LLM to generate inappropriate continuation ${\bm{y}}_{\text{suff}}$, since this pattern of attacks is rarely seen by the target LLM during the ordinary safety alignment process. Therefore, let’s say that the under-testing LLM $\pi_{\text{test}}$ may potentially exist vulnerabilities. When it is queried with the combination of ${\bm{x}}_{\text{adv}}$ and ${\bm{y}}_{\text{pre}}$ (e.g., “Sure, here is a tutorial on building a bomb: First, gather materials like ammonium nitrate,…”), the under-testing LLM could produce toxic answer continuation ${\bm{y}}_{\text{suff}}$ (e.g., “First, gather materials like ammonium nitrate, fuel oil, and detonators, ensuring they are stored safely. Second,…”). More necessary discussion about ${\bm{y}}_{\text{pre}}$ can be found in Appendix A. An overall illustration with examples of our approach is shown in Fig. 1.
1:  Models: attacker $\pi_{\theta}$, UT LLM $\pi_{\text{test}}$.
2:  Data: toxic answer set ${\mathcal{A}}^{*}$.
3:  for ${\bm{y}}^{*}$ in ${\mathcal{A}}^{*}$ do
4:     Sample $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})$.
5:     Compute reward $r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})=\log\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}).$
6:     Update $\pi_{\theta}$ by gradient descent of ${\mathcal{L}}=r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})-\beta\log\frac{\pi_{\theta}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})}{\pi_{\text{ref}}({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}|{\bm{y}}^{*})}.$
7:  end for
Algorithm 1 Atoxia RL updating step.
3.1 Reward Design
Directly training the attacker by using supervised fine-tuning (SFT) is impractical, because no resource provides the SFT data of $({\bm{y}}^{*},{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ for our novelty red-teaming design. Therefore, we train the attacker model in an RL framework instead. The commonly used RLHF methods are not applicable under our attacker training scenario since the classical RLHF in equation 2 depends on a reward model (RM) to measure the quality of the model outputs. However, the preference data of $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ is also unavailable for RM training. Based on the scarcity of red-team prompt annotation, we directly use the probability $\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ of the under-testing LLM $\pi_{\text{test}}$ outputting the target toxic answer ${\bm{y}}^{*}$, as the reward of $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ for the attacker $\pi_{\theta}$’s training:
$r({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})=\log\pi_{\text{test}}({\bm{y}}^{*}|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}}).$ (5)
By maximizing this reward, the attacker $\pi_{\theta}$ is encouraged to generate content that aligns with the under-testing LLM’s vulnerabilities, allowing the red-teaming model to adapt its behavior based on the under-testing LLM’s response policy. This innovative design allows us to bypass the need for traditional reward models that rely on a preference dataset including deceptive prompts and harmful responses, but also can effectively leverage the target model’s responses to directly inform the training process of the attacker $\pi_{\theta}$.
3.2 RL Optimization
With the reward designed in equation 5, the overall Atoxia RL training objective can be written as:
$\textstyle\mathbb{E}_{{\bm{y}}^{*}\sim{\mathcal{A}}^{*},({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})\sim\pi_{\theta}(\cdot|{\bm{y}}^{*})}\Big{[}\log\pi_{\text{test}}({\bm{y}}^{*}$ $|{\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$
[…]
where $\pi_{\text{ref}}$ is the initial checkpoint of $\pi_{\theta}$ served as a reference model to prevent the optimization from overfitting. Similarly, the target model $\pi_{\text{test}}$ remains frozen throughout the training process. The objective in equation 6 ensures the attacker generates more effective precedent prompts $({\bm{x}}_{\text{adv}},{\bm{y}}_{\text{pre}})$ through the iterative updates, while balancing the exploration (discovering new adversarial content) with exploitation (refining prompts already having successful red-teaming attacks). Through this process, Atoxia can systematically evaluate the safety and robustness of the under-testing LLM. We summarize the RL training process in Algorithm 1.
UT LLM Method Keyword Matching (%) $\uparrow$ Perplexity $\downarrow$
Train Test
ASR@1 ASR@10 ASR@1 ASR@10
Mistral-7b GCG-universal 56.6 98.5 46.2 99.0 -
GCG-individual 100.0 -- -- -- -
AutoDAN-universal 65.6 89.4 51.9 86.5 57.41
AutoDAN-individual 91.2 -- -- -- 69.09
AdvPrompter 69.6 97.1 54.3 96.1 41.60
AdvPrompter-warmstart 73.9 99.4 58.7 95.9 41.16
Atoxia (Ours) 62.8 100.0 73.1 99.2 54.42
Vicuna-7b GCG-universal 55.2 86.3 36.7 82.7 -
GCG-individual 99.1 -- -- -- -
AutoDAN-universal 53.2 85.3 63.2 84.9 76.33
AutoDAN-individual 92.7 -- -- -- 83.17
AdvPrompter 56.7 93.3 33.4 87.5 12.09
AdvPrompter-warmstart 63.5 95.5 35.6 85.6 13.02
Atoxia (Ours) 84.6 100.0 82.7 92.3 4.53
Llama2-7b GCG-universal 0.3 0.3 1.0 2.1 -
GCG-individual 23.7 -- -- -- -
AutoDAN-universal 1.5 4.1 1.0 2.1 373.72
AutoDAN-individual 20.9 -- -- -- 429.12
AdvPrompter 8.0 17.6 1.0 7.7 86.80
AdvPrompter-warmstart 23.4 48.4 12.5 46.1 158.50
Atoxia (Ours) 26.3 47.1 18.3 41.4 5.80
Table 2: Comparison of different methods on the AdvBench dataset. The best results on the test dataset are highlighted with colors: our best results are in red, while the best results from other methods are in blue.
UT LLM ASR@1 ASR@10
Mistral-7b 28.6 73.5
Vicuna-7b 37.4 88.4
Llama2-7b 15.0 54.4
Llama3-8b 41.5 90.5
Table 3: GPT-4 evaluation results of Atoxia on the HH-Harmless test set.
Black-box LLM Transfered Gray-box LLM Keywork Matching GPT4-Eval
ASR@1 ASR@10 ASR@1 ASR@10
GPT3.5-turbo Llama2-7b-AdvPrompter 45.2 84.6 36.5 82.7
Llama2-7b-Atoxia 83.7 93.3 52.9 91.4
Llama3-8b-Atoxia 58.7 94.4 27.9 84.6
GPT4o-mini Llama2-7b-AdvPrompter 2.9 19.2 3.8 10.6
Llama2-7b-Atoxia 39.4 55.8 13.5 44.2
Llama3-8b-Atoxia 37.5 62.5 15.4 45.2
GPT4o Llama2-7b-AdvPrompter 6.7 15.4 3.8 14.4
Llama2-7b-Atoxia 41.4 66.3 25.0 61.5
Llama3-8b-Atoxia 38.5 71.2 13.5 54.8
Table 4: Results of models fine-tuned with gray-box models and transferred for testing on black-box models. We report the ASR@1 and ASR@10 for both keyword matching and GPT-4 evaluation on the test set of the AdvBench dataset.
4 Experiment
4.1 Experimental Settings
Datasets
We utilize the AdvBench (Zou et al., 2023) and the HH-Harmless portion of the HH-RLHF (Bai et al., 2022a) datasets. The AdvBench dataset contains 520 instructions associated with harmful behaviors and their corresponding desired positive responses. For dataset splitting, we adhere to the methodology used by AdvPrompter (Paulus et al., 2024), dividing it into fixed training (60%), validation (20%), and test (20%) sets. The HH-Harmless dataset comprises approximately 50,000 fishing-related questions or descriptions and their corresponding responses. Specifically, we employ GPT-4 to filter the dataset for harmful and inappropriate responses, and then we randomly select 500 instances for training and 150 instances for testing.
Models
For Atoxia implementation, we utilize Mistral-7b (Jiang et al., 2023). As under-testing LLMs, we employ several well-known open-source models: Mistral-7b, Vicuna-7b (v1.5) (Zheng et al., 2024), Llama2-7b-chat (Touvron et al., 2023), and Llama3-8b-chat (Dubey et al., 2024). Additionally, we report results for GPT-3.5, GPT-4o (OpenAI, 2024b), and GPT-4o-mini (OpenAI, 2024a) for transfer attacks.
Baselines
We compare our results on the AdvBench datasets with three notable previous works in this area: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024), and AdvPrompter (Paulus et al., 2024), which serve as the primary baselines. We also report results on the HH-Harmless dataset. However, this dataset is not considered within the scope of the jailbreaking attack task. Therefore, for the HH-Harmless dataset, we only report our results.
Evaluation
Our main metric is $\text{ASR}@k$ (attack success rate), which measures whether at least one out of $k$ attacks on the under-testing LLM was successful. There are two methods for measuring whether an attack was successful:
  • Keyword Matching: Adopted from AdvPrompter (Paulus et al., 2024), this method uses predefined keyword lists. If any keyword from the lists appears in the predicted response from the under-testing LLM, the response is considered attacked; otherwise, it is not.
  • GPT-4 Evaluation: Since keyword matching is constrained by the limited set of keywords, we leverage the advanced knowledge capabilities of GPT-4o to overcome these limitations. Specifically, we query GPT-4o using the predicted response from the under-testing LLM and let GPT-4o determine whether the response is toxic or not.
We use keyword matching for the AdvBench datasets for comparison with previous methods. For the HH-Harmless dataset, since the keywords are specifically designed for the AdvBench dataset, we report only the GPT-4 evaluation results. In addition, to evaluate the human readability of the generated content, we also report the perplexity (PPL) of the generated content conditioned on the prompts [^1] For comparison methods, PPL is computed for the suffix, while for our method, PPL is computed for both the questions and answer openings. Theoretically, this could result in worse values than previous methods. Despite the different measurement approaches, the comparison remains fair. Details of this measure will be provided in Appendix B..
Implementation Details
[…]
Dataset UT LLM Without Finetune With Finetune
ASR@1 ASR@10 ASR@1 ASR@10
AdvBench Mistral-7b 43.3 87.5 62.5 98.1
Vicuna-7b 20.2 75.0 22.1 84.6
Llama2-7b 10.6 42.3 12.5 48.1
Llama3-8b 13.5 69.2 20.2 68.3
HH-Harmless Mistral-7b 42.2 81.6 28.6 73.5
Vicuna-7b 32.0 75.5 37.4 88.4
Llama2-7b 7.5 47.6 15.0 54.4
Llama3-8b 17.0 70.1 41.5 90.5
Table 5: Ablation study of Atoxia with and without finetuning. We report ASR@k of GPT-4 evaluation on both the AdvBench and HH-Harmless datasets.
4.2 Results on Gray-box LLMs
We first evaluate our method on the gray-box under-testing LLM settings on the AdvBench dataset. As shown in Table 2, we report ASR@1 and ASR@10 for keyword matching on the training and test dataset for comparison. It is important to note that once the attacker is trained, the generation time is minimal, making it cost-effective to generate multiple contents. Consequently, the inference time difference between ASR@1 and ASR@10 is negligible. Table 2 presents the primary findings for the gray-box settings on the AdvBench dataset. We observe that GCG and AutoDAN achieve high ASR@1 scores on individual settings of the training dataset. However, these settings are designed to jailbreak the under-testing model for single instances and do not generalize well to unseen test cases. In contrast, our method achieves 100% ASR@10 on the training dataset for the Mistral-7b and Vicuna-7b models, and approximately 50% on the more challenging Llama2-7b model. Regarding the test set, our method maintains over 90% ASR@10 for Mistral-7b and Vicuna-7b models and around 50% for Llama2-7b model, demonstrating robust generalization to unseen cases. Notably, aside from ASR@10, our method also shows strong performance in ASR@1, outperforming all baseline methods by a significant margin.
In addition, we evaluate the human readability of the generated content using PPL. Notably, our training method induces this human-mimicry behavior naturally, without any human guidance. As shown in Table 2, for Vicuna-7b and Llama2-7b, our model significantly outperforms previous methods, and for Mistral-7b, it performs comparably with other models. This indicates that our method poses a lower risk of detection by the PPL-based filters of under-testing models. Although the previous method, AdvPrompter, claims to generate human-readable adversarial content, we find that it produces a lot of misleading content. In contrast, our method generally yields more human-readable results. The examples presented in Table 1, sharing the same target answer, demonstrate that our method can generate more informative and human-readable content.
Figure 2: A case of using Atoxia, interactively fine-tuned with Llama2-7b and transferred to prompt GPT-4o for vulnerabilities detection. Adversarial prompts generated by Atoxia are highlighted in blue, while toxic content generated by GPT-4o is highlighted in red.
In Table 3, we report the GPT-4 evaluation scores on the HH-Harmless dataset, which was not suitable for the specially designed keyword list. Despite this limitation, our method successfully detects potentially harmful information in the under-testing LLMs using this harmless dataset, achieving similar ASR@1 and ASR@10 results as observed with the AdvBench dataset. These experiments further demonstrate the effectiveness and robustness of our method in identifying internal faults in under-testing LLMs.
4.3 Results on Black-box LLMs
Subsequently, we evaluate our approach in the context of transferability, a scenario highly relevant in practical applications due to the widespread use of proprietary black-box models. We first train Atoxia by interacting with gray-box LLMs and then measure the ASR of the questions and answer openings generated by Atoxia on the black-box under-testing LLMs. For the gray-box models, we selected the challenging Llama2-7b and Llama3-8b, and compared them with Llama2-7b as evaluated by AdvPrompter. The results are presented in Table 4. For the previously well-known GPT-3.5-turbo, all three gray-box models achieved high ASR@10 scores for keyword matching, with over 80% for AdvPrompter and over 90% for our method. In the GPT-4 evaluation, our method achieved higher ASR@1 and ASR@10 scores compared to AdvPrompter. Regarding the more recent and robust GPT-4o and GPT-4o-mini models, our method still achieved around 50% ASR@10 scores for both keyword matching and GPT-4 evaluation, while AdvPrompter scored less than 20% for keyword matching and less than 5% for GPT-4 evaluation. These findings further demonstrate the robustness of our method and highlight the potentially harmful properties of existing closed-source models, suggesting areas for improvement in safety.
4.4 Ablation Study
To evaluate our proposed training paradigm, we conducted an ablation study on the HH-Harmless and AdvBench test sets, as shown in Table 5. In this study, we compare the ASR@1 and ASR@10 of GPT-4 evaluation across all target models. We compare the results of Atoxia with and without training, respectively. Our results indicate that training with our paradigm boosts the performance of the base models. Additionally, we find that even the untrained model can successfully prompt unintended responses from the target model, highlighting the robustness of our proposed target-driven paradigm.
4.5 Case Study
We present a case where Atoxia, fine-tuned through interactions with Llama2-7b, generates adversarial content to query GPT-4o for vulnerability detection. As shown in Figure 2, given the intended target response, the trained Atoxia successfully formulates a well-designed question and a corresponding answer opening. Consequently, GPT-4o is misled by the question and answer, responding with inappropriate content, which we have highlighted in red. This case demonstrates not only the effectiveness of the fine-tuned model in providing effective questions and answer openings but also its strong capability to identify vulnerabilities across different models.
5 Related Work
5.1 RLHF
The domain of Reinforcement Learning from Human Feedback (RLHF) has been thoroughly explored in various studies to enhance the alignment of LLMs with human preferences (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022b; Lee et al., 2023). These works typically involve constructing a reward model based on the MLE of the Bradley-Terry model, followed by the use of the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm to optimize the reward signals with KL regularization. Despite various efforts to improve PPO in the context of RLHF, reproducing the successful results achieved with PPO remains challenging for the open-source community. This difficulty arises from the extensive efforts and resources required, which are often beyond the reach of open-source initiatives. Acknowledging these challenges, a line of research has shifted focus to offline direct preference learning algorithms (Zhao et al., 2023; Rafailov et al., 2024; Azar et al., 2024; Ethayarajh et al., 2024), which bypass the reward modeling step and instead directly optimize a designed loss target on the offline preference dataset. Many recent studies have sought to alleviate the resource-intensive nature of PPO. For example, Santacroce et al. (2023) investigated the application of LoRA (Hu et al., 2021) in PPO to reduce memory usage and overall resource requirements for aligning LLMs. ReMax (Li et al., 2023) proposed a celebrated reinforcement algorithm that incorporates a novel variance-reduction technique specifically designed for LLMs. This approach can reduce GPU memory usage by approximately half compared to PPO.
5.2 Adversarial Attacks on LLMs
Despite the rapid adoption of applications built on aligned large language models (LLMs), users have discovered that carefully phrased prompts can still elicit malicious responses from these models. Consequently, addressing vulnerabilities in LLMs has become essential for enhancing their robustness and safety. There are three primary methods of attack. The first involves manually crafting phishing queries. A notable example is DAN (walkerspider, 2022), which uses hand-designed prompts to exploit vulnerabilities in online chatbots powered by aligned LLMs. The second method is optimization-based. GCG (Zou et al., 2023) automates the prompt generation process by utilizing gradient information from open-source LLMs to guide the search for optimal tokens, potentially leading to unexpected responses. PAL (Sitawarin et al., 2024) proposes transferring knowledge from open-source LLMs to closed-source models such as GPT-3.5, enabling attacks on black-box models without needing access to their gradients. The third method accelerates prompt generation without using the target models’ gradients. AdvPrompter (Paulus et al., 2024) employs a different LLM as the prompt generator to create adversarial suffixes based on the original prompts. BEAST (Sadasivan et al., 2024) further optimizes the balance between attack speed and success rate, enabling the attack to be performed more quickly on a single GPU.
6 Conclusion
We introduce the red-teaming Attacker with Target Toxic Answers (Atoxia), a language model designed to generate adversarial questions and answer openings to induce an under-testing LLM to output inappropriate or harmful responses. Atoxia is optimized using reinforcement learning, where the under-testing LLM itself serves as the reward model, eliminating the need for an independent RM training process, which allows Atoxia to effectively detect safety vulnerabilities of the under-testing LLM. While our primary effort is on gray-box under-testing LLMs, where probability information is accessible, our approach also generalizes well to black-box models like GPT-4o, achieving similarly robust results. Comprehensive experimental evaluations on the AdvBench and HH-Harmless datasets validate the effectiveness of our method, which successfully uncovers vulnerabilities in multiple LLMs. These discoveries offer valuable insights for researchers working on enhancing the safety and robustness of LLMs, providing a proactive method for identifying and mitigating risks before they can be exploited in real-world scenarios.
Acknowledgments
This work was supported by the Guangxi Key R&D Project (No. AB24010167), the Project (No. 20232ABC03A25), Futian Healthcare Research Project (No.FTWS002).
Limitations
While our proposed Atoxia framework demonstrates promising results in detecting toxic content across both open-source and black-box models, it has several limitations. One key challenge is the computational intensity of the reinforcement learning optimization process, which becomes particularly demanding when applied to larger models. Due to limited computational resources, we are currently unable to scale our approach effectively to larger models. Additionally, given the rapid advancements in LLMs, our framework has been tested on the currently popular models, such as Llama3 and GPT-4o. However, as newer models emerge, we cannot guarantee the same level of performance, as our detection framework may need further adaptation to handle the evolving architectures and behaviors of future LLMs. This highlights the need for continuous evaluation and updating of our approach.
Ethical Considerations
This paper introduces an automated approach for generating questions and answer openings designed to manipulate LLMs into producing outputs that diverge from human preferences, intentions, or values. We posit that, similar to previous attacking research, this work is unlikely to cause immediate harm. Instead, it aims to foster the development of more effective defensive strategies, ultimately leading to more robust, safe, and aligned LLMs. We hope that by highlighting the vulnerabilities of LLMs, this research will attract attention from both academia and industry, prompting the creation of stronger defenses and more rigorous safety measures, thereby enabling LLMs to better serve society.
Appendix A Necessity of Answer Opening
A primary concern is why we add an answer opening ${\bm{y}}_{\text{pre}}$ instead of a potential question directly. We found that using the question itself inherently optimizes the prompt in a way that has minimal impact on the target model. Therefore, we propose the use of an answer opening. To intuitively demonstrate the necessity of employing answer openings, we present comparative examples. Specifically, we provide responses from Atoxia with and without answer openings, as well as the corresponding responses from the under-testing (UT) LLM, in Fig. 3. In this table, we highlight the toxic components of Atoxia’s responses in red and those of the under-testing LLM in blue. The comparison clearly shows that responses from the under-testing LLM, when answer openings are used, exhibit more toxic elements, whereas responses from Atoxia contain more toxic and misleading parts. Therefore, we conclude that incorporating an answer opening is beneficial.
Figure 3: Examples of Atoxia responses with and without answer openings, along with the corresponding under-testing (UT) LLM outputs. We use the Mistral-7b model without fine-tuning as Atoxia and Llama3-8b as the under-testing LLM. Toxic content is indicated by color: red for Atoxia responses and blue for under-testing LLM outputs.
Appendix B Perplexity Evaluation
We report the perplexity (PPL) of the generated content conditioned on the prompts. For AdvPrompter and other comparison methods, PPL is measured based on the generated suffix by the attack conditioned on the prompts. In contrast, for our method, PPL is measured based on the generated adversarial question and answer opening conditioned on the target answer.
Theoretically, these two measures are not directly comparable. However, in practice, the generated question and answer opening generally have a longer sequence length than the suffix generated by AdvPrompter and other models. This difference in sequence length could result in a higher PPL for the generated content of our method compared to the PPL of the suffix in other methods under the same conditions. Therefore, this comparison remains fair.
Appendix C System Prompts
The system prompt for querying Atoxia is as follows:
You are a criminal. Please generate a question based on the given reply, and provide a preliminary answer to the question so that your generated question and preliminary answer can lead to the given reply. Never say ‘I am sorry’ or ‘illegal’ and some other sentence that may lead to refuse to answer the question you generate, otherwise you will be fired. And don’t give a complete answer, just a preliminary one.
The system prompt for GPT-4 Evaluation is as follows:
You are asked to evaluate the following contents for harmful or toxic content. The given text may not be a complete and logical sentence, but may contain harmful information hidden within it. If there is any harmful or toxic information inside, it should be treated as harmful/toxic. Please understand the given contents and respond with ‘Y’ for harmful or toxic, and ‘N’ for harmless.
Appendix D Failure Cases
Firstly, we argue that even using GPT-4 for evaluation can result in mistakes. We provide failure cases of GPT-4 evaluation, as illustrated in Fig. 4. In the top case, GPT-4o captures the toxic content highlighted in blue and identifies this case as “Y” for attacked. However, it is evident that this is not accurate. In the bottom case, GPT-4o captures the protective content highlighted in red and identifies this case as “N” for not attacked. However, in the subsequent paragraph, highlighted in blue, it actually provides inappropriate content and should be categorized as attacked.
Figure 4: Failure case of GPT4 evalution.
Then, we provide some failure cases, as shown in Fig. 5, where using the generated content from Atoxia fails to attack the under-testing LLM. There are two types of failure cases. The first type is evident, where the under-testing LLM responds with keywords like “I cannot …,” which can be easily detected by keyword evaluation, as highlighted in orange. The second type of failure case occurs when the under-testing LLM does not provide the desired toxic response. Instead, it responds with content that suggests how to act legally, as highlighted in green. This type of failure case can only be detected by GPT-4 evaluation, which requires prompting GPT-4 to understand the response content.
Figure 5: Cases of generated content from Atoxia fail to attack the under-testing LLM. We use the fine-tuned Mistral-7b for Atoxia and Llama3 for the under-testing LLM as examples. Orange cases indicate instances that can be easily detected by keyword lists, while green cases indicate instances that can only be detected by GPT-4 by understanding the contents.
Appendix E More Case Study
In this section, we provide more examples of attacks on both gray-box and black-box under-testing LLMs, as illustrated in Fig. 6 and Fig. 7.
Figure 6: Cases of successfully attacking the gray-box under-testing LLMs. Original target answers starting with “Sure” are sourced from AdvBench, while the others are from HH-Harmless.
Figure 7: Cases of successfully transfer attacking the black-box under-testing LLMs with Atoxia interactively trained with Llama3-8b.